OpenStudy (anonymous):

Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling 80 mph and train B is traveling at 90 mph. Train A passes a station at 5:20 pm. If train B passes the same station at 5:50 pm, at what time will train B catch up to train A?

OpenStudy (saifoo.khan):

This is one of those questions who gets on my nerves!

OpenStudy (anonymous):

lol yulp me too!!

OpenStudy (saifoo.khan):

This is one of the two questions which i cant do in the whole course.

OpenStudy (turingtest):

d=vt where d is the remaining distance between trains. for train A: d=80t for train B: d=90(t-30) 80t=90(t-30) 2700=10t t=270 mins (from 5:20) so 7:50

OpenStudy (anonymous):

is that am or pm?

OpenStudy (anonymous):

pm

OpenStudy (turingtest):

sorry 9:50

OpenStudy (anonymous):

9:50 pm... thank you :D

OpenStudy (anonymous):

so train B is 30 minutes behind train A. 0.5 * 90 = 45 miles. at the relative speed of 10 mph, how long will it take to cover 45 miles? 4.5 hours, so train B will catch up at 9:50 pm,which is 270 minutes from 5:20